OpenStudy (anonymous):

Calculate the velocity of a 1650 kg satellite that is in a circular orbit of 4.2 x 106m above the surface of a planet, which has a radius 3.95 x 105 m and the period of the orbit is 2.0 hours.

OpenStudy (anonymous):

@YanaSidlinskiy one more..? this one has killed me for the last 2 days

OpenStudy (anonymous):

@chmvijay

OpenStudy (yanasidlinskiy):

Ok. Sorry. I made you wait..

OpenStudy (anonymous):

No its fine

OpenStudy (yanasidlinskiy):

if a satellite is moving in a orbit of radius R with velocity V and time period T then its time period is given by \(T=\huge\frac{ 2\pi R }{ V}\) so velocity will be \[V = \frac{ 2\pi R }{ T }\]

OpenStudy (yanasidlinskiy):

What did you get?

OpenStudy (anonymous):

So 2*pi*3.95*10^5/2 right?

OpenStudy (anonymous):

I FINALLY GOT IT 1.24*10^6m/s

OpenStudy (anonymous):

Please tell me thats right

OpenStudy (yanasidlinskiy):

Yep!:) You are correct! Sorry I had to run an errand..

OpenStudy (anonymous):

Its fine and thank you sooooooooooo much!!!!!!

OpenStudy (yanasidlinskiy):

You're Welcome!! :D